# SpGesture: Source-Free Domain-adaptive sEMG-based Gesture Recognition with Jaccard Attentive Spiking Neural Network

### v1

@@ Abstract
Surface electromyography (sEMG) based gesture recognition offers a natural and intuitive interaction modality for wearable devices. Despite significant advancements in sEMG-based gesture-recognition models, existing methods often suffer from high computational latency and increased energy consumption. Additionally, the inherent instability of sEMG signals, combined with their sensitivity to distribution shifts in real-world settings, compromises model robustness. To tackle these challenges, we propose a novel SpGesture framework based on Spiking Neural Networks, which possesses several unique merits compared with existing methods: (1) Robustness: By utilizing membrane potential as a memory list, we pioneer the introduction of Source-Free Domain Adaptation into SNN for the first time. This enables SpGesture to mitigate the accuracy degradation caused by distribution shifts. (2) High Accuracy: With a novel Spiking Jaccard Attention, SpGesture enhances the SNNs’ ability to represent sEMG features, leading to a notable rise in system accuracy. To validate SpGesture’s performance, we collected a new sEMG gesture dataset which has different forearm postures, where SpGesture achieved the highest accuracy among the baselines ($89.26\%$). Moreover, the actual deployment on the CPU demonstrated a system latency below 100ms, well within real-time requirements. This impressive performance showcases SpGesture’s potential to enhance the applicability of sEMG in real-world scenarios. We will release our code and model to facilitate further research.
footnotetext: Work done during an internship at Guangzhou MetaHCI, Co., Ltd.footnotetext: Corresponding Author.
@@ 1 Introduction
Although SNNs are computationally efficient, they struggle to match the accuracy of ANN-based models [15]. In particular, the major problem is that the binary and sparse feature representations make it difficult to perform regular contiguous similarity computations. This limitation hinders expressive operations like attention mechanisms and advanced representation alignment algorithms, such as domain adaptation. For example, attention-based structures like the Transformer have demonstrated remarkable performance in both Natural Language Processing [16, 40, 8] and Computer Vision [26, 30, 39] tasks, leading to a wave of attention-centric architecture designs, underscoring the importance and versatility of attention mechanisms in deep learning. However, with the proportion of ‘1’s typically less than $5\%$, the dot product in cosine similarity inherent to attention mechanisms tends to yield results close to zero [65]. Existing work often first converts spike signals into continuous values for similarity calculations, but this can increase the inference latency and energy consumption of SNNs. There is still a lack of methods for directly implementing advanced operations on spike features in SNN for realizing effective sEMG-based gesture recognition systems.
@@ 4.1.2 Spiking Jaccard Attention
$\mathrm{SJA}\left(\mathbf{Q},\mathbf{K}\right)=\frac{\sum_{ij}\min\left(q_{ij},k_{ij}\right)}{\sum_{ij}\max\left(q_{ij},k_{ij}\right)+\epsilon}\ \mathbf{V},$ (3)
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ represent the query, key, and value matrices, respectively, and $q_{ij}$ and $k_{ij}$ are the corresponding elements in the query and key matrices.
@@ 4.2 Spiking Source-Free Domain Adaptation based on Membrane Potential Memory
$\mathcal{S}_{k}$ =$=\begin{dcases}\text{mode}\left(\text{argmax}\left(\left\{\mathcal{M}\right\}_{1}^{k}\right)\right),&\text{with probability }p,\\ \text{random}\left(\text{argmax}\left(\left\{\mathcal{M}\right\}_{1}^{k}\right)\right),&\text{with probability }(1-p),\end{dcases}$ (6)
@@ A.1 Details of Data Collection
The experiments were carried out using the DataLITE wireless LE230 and DataLITE PIONEER, commercial sEMG acquisition systems from Biometrics Ltd.https://www.biometricsltd.com/ The device’s sampling rate is 2000Hz, allowing for high-resolution data capture of the electrical activities in the muscles during the performance of gestures. Eight LE230 sEMG sensors were uniformly and equidistantly affixed to the surface of the participant’s right forearm.

### v2

@@ Abstract
Surface electromyography (sEMG) based gesture recognition offers a natural and intuitive interaction modality for wearable devices. Despite significant advancements in sEMG-based gesture-recognition models, existing methods often suffer from high computational latency and increased energy consumption. Additionally, the inherent instability of sEMG signals, combined with their sensitivity to distribution shifts in real-world settings, compromises model robustness. To tackle these challenges, we propose a novel SpGesture framework based on Spiking Neural Networks, which possesses several unique merits compared with existing methods: (1) Robustness: By utilizing membrane potential as a memory list, we pioneer the introduction of Source-Free Domain Adaptation into SNN for the first time. This enables SpGesture to mitigate the accuracy degradation caused by distribution shifts. (2) High Accuracy: With a novel Spiking Jaccard Attention, SpGesture enhances the SNNs’ ability to represent sEMG features, leading to a notable rise in system accuracy. To validate SpGesture’s performance, we collected a new sEMG gesture dataset which has different forearm postures, where SpGesture achieved the highest accuracy among the baselines ($89.26\%$). Moreover, the actual deployment on the CPU demonstrated a system latency below 100ms, well within real-time requirements. This impressive performance showcases SpGesture’s potential to enhance the applicability of sEMG in real-world scenarios. The code is available at https://github.com/guoweiyu/SpGesture/.
@@ 1 Introduction
Although SNNs are computationally efficient, they struggle to match the accuracy of ANN-based models [15]. In particular, the major problem is that the binary and sparse feature representations make it difficult to perform regular contiguous similarity computations. This limitation hinders expressive operations like attention mechanisms and advanced representation alignment algorithms, such as domain adaptation. For example, attention-based structures like the Transformer have demonstrated remarkable performance in Natural Language Processing [16, 40, 8, 69], Computer Vision [26, 30, 39, 70, 68], Time-Series Processing [76, 77, 75] and Decision Making tasks[56, 57], leading to a wave of attention-centric architecture designs, underscoring the importance and versatility of attention mechanisms in deep learning. However, with the proportion of ‘1’s typically less than $5\%$, the dot product in cosine similarity inherent to attention mechanisms tends to yield results close to zero [67]. Existing work often first converts spike signals into continuous values for similarity calculations, but this can increase the inference latency and energy consumption of SNNs. There is still a lack of methods for directly implementing advanced operations on spike features in SNN for realizing effective sEMG-based gesture recognition systems.
@@ 4.1.2 Spiking Jaccard Attention
$\mathrm{SJA}\left(\mathbf{Q},\mathbf{K}\right)=\frac{\sum_{i}\min\left(q_{i},k_{i}\right)}{\sum_{i}\max\left(q_{i},k_{i}\right)+\epsilon}\ \mathbf{V},$ (3)
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ represent the query, key, and value matrices, respectively, and $q_{i}$ and $k_{i}$ are the corresponding elements in the query and key matrices.
First, we consider the channel-wise uniform weighting method. This approach implies that the same weighting coefficient is applied to all elements along the channel dimension of $\mathbf{V}$. In this case, the attention weight is computed as a scalar, calculated by aggregating the elements within each channel: where $i$ is the index of the elements within the channel. The resulting scalar is then used as a weighting factor applied to each channel of $\mathbf{V}$:
$\mathbf{V}_{\text{new}}[:,c,:]=\mathrm{Jaccard}(\mathbf{Q},\mathbf{K})\cdot\mathbf{V}[:,c,:],$ (4)
where $c$ represents the channel index. In this way, the values across all channels are scaled by the same weighting factor, thereby maintaining consistency across different channels.
Second, we consider the element-wise weighting method. In this case, the $\mathrm{Jaccard}(\mathbf{Q},\mathbf{K})$ result is computed independently for each element position $q,k$. This means that the attention weight for each element is obtained by calculating the value for the corresponding elements in $\mathbf{Q}$ and $\mathbf{K}$ at that position. These weights are then applied element-wise to $\mathbf{V}$:
$\mathbf{V}_{\text{new}}[:,c,n]=\mathrm{Jaccard}(\mathbf{Q},\mathbf{K})[:,c,n]\cdot\mathbf{V}[:,c,n],$ (5)
where $n$ represents the index along the sequence length. Under this element-wise weighting strategy, different positions within $\mathbf{V}$ are scaled independently based on their respective attention weights, which enables the model to capture finer-grained features.
The results presented in this paper are derived using the channel-wise weighting approach, as it is more suitable for the characteristics of sEMG data, and we did not validate the element-wise weighting approach due to these characteristics.
These two weighting strategies each have their respective applications: channel-wise uniform weighting is more appropriate for preserving feature consistency, while element-wise weighting is better suited for capturing localized differences. Depending on the computational complexity and the task requirements, an appropriate weighting strategy can be selected to achieve a balance between efficiency and performance.
@@ 4.2 Spiking Source-Free Domain Adaptation based on Membrane Potential Memory
$\mathcal{S}_{k}$ =$=\begin{dcases}\text{mode}\left(\text{argmax}\left(\left\{\mathcal{M}\right\}_{1}^{k}\right)\right),&\text{with probability }(1-p),\\ \text{random}\left(\text{argmax}\left(\left\{\mathcal{M}\right\}_{1}^{k}\right)\right),&\text{with probability }p,\end{dcases}$ (8)
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
This work is partly supported by the National Key Research and Development Program of China (No. 2023YFF0725000), the National Natural Science Foundation of China (No. 92370204, 62306255), the Natural Science Foundation of Guangdong Province (No. 2024A1515011839), the Guangzhou-HKUST(GZ) Joint Funding Program (No.2023A03J0008), and the Education Bureau of Guangzhou Municipality. Yijie Xu acknowledges the support from the modern matter laboratory, HKUST(GZ).
@@ A.1 Details of Data Collection
The experiments were carried out using the DataLITE wireless LE230 and DataLITE PIONEER, commercial sEMG acquisition systems from Biometrics Ltd. [^1] https://www.biometricsltd.com/ The device’s sampling rate is 2000Hz, allowing for high-resolution data capture of the electrical activities in the muscles during the performance of gestures. Eight LE230 sEMG sensors were uniformly and equidistantly affixed to the surface of the participant’s right forearm.